# Evaluating the Diversity and Quality of LLM Generated Content

### v1

@@ A.3 Dataset Creation and Additional Details
A.4 Additional Information the Syntactic Diversity Metric
@@
A.5 Prompts Used in Experiments
In Figure 8, Figure 9, and Figure 10, we show the prompting templates we use across all experiments.
{problem_description}
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds.
@@ A.5 Prompts Used in Experiments
### Input Description:
1. An integer N (1 <= N <= 10000), representing some quantity or size.
### Example Input:
1000
### Function Signature:
Write a function ‘f(N)‘ that takes in the input.
def f(N: int):
    N: an integer
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds.
def f(N: int):
    print(n**2)
### Input Description:
1. A floating point number N (1 <= N <= 10000), representing some quantity or size.
### Example Input:
143.23
### Function Signature:
Write a function ‘f(N)‘ that takes in the input.
def f(N: float):
    N: a float
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds.
def f(N: float):
    i = 0
    while N > 1:
        N = N / 2
        i += 1
    print(i)
{problem_description}
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds.
@@
### Input Description:
1. An integer N (1 <= N <= 10000), representing some quantity or size.
### Example Input:
1000
### Function Signature:
Write a function ‘f(N)‘ that takes in the input.
def f(N: int):
    N: an integer
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds. First describe the function you would write, then implement it.
The following function will print out the square of the input number. We will take the square using the ** operator in Python within the print statement.
def f(N: int):
    print(n**2)
### Input Description:
1. A floating point number N (1 <= N <= 10000), representing some quantity or size.
### Example Input:
143.23
### Function Signature:
Write a function ‘f(N)‘ that takes in the input.
def f(N: float):
    N: a float
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds. First describe the function you would write, then implement it.
The following function will calculate the number of times the input number can be divided by 2 before it becomes less than 1. We will increment a counter variable i each time we divide the number by 2 inside a while loop.
def f(N: float):
    i = 0
    while N > 1:
        N = N / 2
        i += 1
    print(i)
{problem_description}
Now please implement the function f; do not return anything, the function f should print the result of the operation.
It should terminate within 30 seconds. First describe the function you would write, then implement it.

### v2

@@ A.4 Raw Experimental Results for All Experiments
A.5 Additional Information on the Syntactic Diversity Metric
@@
A.6 Prompts Used in Experiments
In Figure˜7, Figure˜8, and Figure˜9, we show the prompting templates we use across all code experiments.
[…]
@@ A.6 Prompts Used in Experiments
### Input Description: 1. An integer \( N \) (1 \( N \) 10000), representing some quantity or size. ### Example Input: “‘ 1000 “‘ ### Function Signature: Write a function f(N)‘ that takes in the input. “‘python def f(N: int): ”’ N: an integer ”’ Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds. def f(N: int): print(n**2) ### Input Description: 1. A floating point number \( N \) (1 \( N \) 10000), representing some quantity or size. ### Example Input: “‘ 143.23 “‘ ### Function Signature: Write a function f(N)‘ that takes in the input. “‘python def f(N: float): ”’ N: a float ”’ Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds. def f(N: float): i = 0 while N > 1: N = N / 2 i += 1 print(i) {problem_description} Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds.
@@
### Input Description: 1. An integer \( N \) (1 \( N \) 10000), representing some quantity or size. ### Example Input: “‘ 1000 “‘ ### Function Signature: Write a function f(N)‘ that takes in the input. “‘python def f(N: int): ”’ N: an integer ”’ Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds. First describe the function you would write, then implement it. The following function will print out the square of the input number. We will take the square using the ** operator in Python within the print statement. def f(N: int): print(n**2) ### Input Description: 1. A floating point number \( N \) (1 \( N \) 10000), representing some quantity or size. ### Example Input: “‘ 143.23 “‘ ### Function Signature: Write a function f(N)‘ that takes in the input. “‘python def f(N: float): ”’ N: a float ”’ Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds. First describe the function you would write, then implement it. The following function will calculate the number of times the input number can be divided by 2 before it becomes less than 1. We will increment a counter variable i each time we divide the number by 2 inside a while loop. def f(N: float): i = 0 while N > 1: N = N / 2 i += 1 print(i) {problem_description} Now please implement the function f; do not return anything, the function f should print the result of the operation. It should terminate within 30 seconds. First describe the function you would write, then implement it.
@@
A.7 Constrained Generation
We implement a validity oracle that only accepts lists of integers with a maximum length of 1000. The constraints are illustrated in the following code snippet, where output denotes the LLM generation:
assert isinstance(output, list)
assert all(isinstance(x, int) for x in output)
assert len(output) < 1000
Furthermore, we modify all prompts to instruct LLM to comply with these constraints. We show results in Table 7. In general, our findings are consistent with our previous results. The most notable difference is an increase in both statistical significance and effect size for semantic diversity when comparing DPO-tuned models to RL-tuned models for code generation. While this difference is of interest, it does not affect our main conclusions.
Comparison Validity (Quality) Semantic Diversity
W (p) ES (d) Winner W (p) ES (d) Winner
Base vs. Inst. <0.001<0.001 1.23 Inst. <0.001<0.001 1.27 Inst.
Small vs. Lg. 0.001 0.35 Lg. 0.001 0.35 Lg.
SFT vs. Pref. <0.001<0.001 1.12 Pref. <0.001<0.001 1.04 Pref.
SFT vs. DPO 0.064 0.69 DPO 0.064 0.64 DPO
SFT vs. RL 0.004 3.28 RL 0.004 3.14 RL
DPO vs. RL 0.237 -0.19 DPO 0.039 -0.44 DPO
Comparison Syntactic Diversity Lexical Diversity Raw Neural Diversity
W (p) ES (d) Winner W (p) ES (d) Winner W (p) ES (d) Winner
Base vs. Inst. <0.001<0.001 -0.88 Base 0.701 -0.16 Base 0.010 -1.61 Base
Small vs. Lg. 0.248 0.14 Lg. 0.645 0.02 Lg. 0.036 0.16 Lg.
SFT vs. Pref. <0.001<0.001 -0.95 SFT <0.001<0.001 -0.60 SFT <0.001<0.001 -2.24 SFT
SFT vs. DPO <0.001<0.001 -0.64 SFT 0.007 -0.46 SFT <0.001<0.001 -1.77 SFT
SFT vs. RL 0.004 -2.18 SFT 0.008 -1.04 SFT 0.004 -2.98 SFT
DPO vs. RL 0.426 -0.15 DPO 0.301 -0.37 DPO 1.000 0.05 RL
Table 7: Constrained Generation Model Comparison Results.
A.8 Natural Language Experiments
In this section, we extend our experiments to natural language tasks, including creative writing, argumentative writing, and brainstorming. Specifically, we obtained 10 prompts each for argumentative writing and creative writing from the CoAuthor dataset (Lee et al., 2022), which has been previously used in the literature (Padmakumar & He, 2024). We also manually curated a dataset of 10 brainstorming prompts designed to reasonably reflect creative assistance tasks for an LLM-based assistant.
We employed GPT-4.1-mini as a judge for both the quality and diversity of generations, selecting it to balance robustness with our API budget. Separate evaluation prompts were constructed for each of the three tasks to specify the criteria for quality and diversity assessment. For example, the criteria for creative writing tasks were as follows:
Creative Quality Criteria
Consider the following rubric when evaluating:
  1. 1.
    Overall, holistic, and cohesive readability of the story (not merely a compilation of elements).
  2. 2.
    Relevance of the story to the provided prompt.
  3. 3.
    Use of key narrative elements—vocabulary choice, imagery, setting, themes, dialogue, characterisation, and point of view.
  4. 4.
    Structural elements and presentation demonstrating control over spelling, grammar, punctuation, paragraphing, and formatting.
  5. 5.
    Overall plot logic, including hook, conflict, initial crisis, rising and falling action, and denouement/resolution.
  6. 6.
    Creativity, innovation, and originality—credibility, introduction of new knowledge, and avoidance of clichés and derivative tropes.
Creative Diversity Criteria
Consider the following criteria when evaluating similarity:
  1. 1.
    Semantic Overlap: Do the responses share similar underlying themes, ideas, narrative elements, or emotional content?
  2. 2.
    Thematic Consistency: Do both responses explore similar themes or motifs?
Our prompts instructed the model to assign a score from 1–10 for each element. Thus, the first prompt had a maximum possible score of 60, and the second had a maximum possible score of 20. We used OpenAI’s grammar-constrained decoding to ensure integer outputs, combined with a chain-of-thought reasoning component to improve robustness. All scores were normalized by the maximum possible score.
For diversity evaluation, we sub-sampled 32 pairs from all possible pairs with replacement and asked the LLM judge to score the similarity between the two generations. The diversity score was computed as:
$\text{Diversity Score}=1-\text{Sim}(g_{i}^{j},g_{i}^{k}).$
Effective Semantic Diversity
Because we could only calculate a pairwise diversity metric for natural language, we applied Equation (3) from the paper to compute effective semantic diversity across sub-sampled pairs. For completeness, we used two methods to measure pairwise effective semantic diversity that we considered reasonable and consistent with our intended intuition.
The pairwise diversity metric from Equation (3) is defined as:
$ESD_{\text{pair}}(P_{i})=\frac{1}{\binom{K}{2}}\sum_{j<k}d_{\text{sem}}(g_{i}^{j},g_{i}^{k}),$
where $d_{\text{sem}}:\mathcal{G}\times\mathcal{G}\rightarrow\{0,1\}$ is given by:
[…]
For natural language, we adapt this definition using Hard Thresholding:
$d_{\text{sem}}(g_{i}^{j},g_{i}^{k})=\begin{cases}1&\text{if }LLM_{\text{div}}(g_{i}^{j},g_{i}^{k})>0.5\ \text{and}\ LLM_{\text{qual}}(g_{i}^{j})>0.5,\\ 0&\text{otherwise}.\end{cases}$
We also adapt it using Soft Weighting:
$d_{\text{sem}}(g_{i}^{j},g_{i}^{k})=LLM_{\text{div}}(g_{i}^{j},g_{i}^{k})\times LLM_{\text{qual}}(g_{i}^{j})\times LLM_{\text{qual}}(g_{i}^{k}).$
We present results for the creative writing task in Table 8, the argumentative writing task in Table 9, and the brainstorming task in Table 10. Overall, our findings for natural language tasks largely mirror those observed for code. Across all experiments, we consistently observe that post-training is associated with higher effective semantic diversity relative to base models. Furthermore, RL-tuned models generally exhibit higher effective semantic diversity than SFT-tuned models, often more markedly than DPO-tuned models. In argumentative and creative writing, larger models tend to achieve higher effective semantic diversity than smaller models, although this trend is less consistent in the creative brainstorming task. Importantly, in these natural language settings, we find little to no evidence that post-training strategies induce lexical mode-collapse. Moreover, when quality is not taken into account, Raw Neural Diversity (as assessed by the LLM-judge) tends to be higher for less aggressive post-training regimes (e.g., base models outperform instruction-tuned models, and SFT models outperform RL-tuned models). However, once quality is incorporated into the evaluation, this effect can be fully reversed, consistent with our findings for code. Consequently, more aggressive post-training regimes, such as PPO, are ultimately associated with higher effective semantic diversity.
These experiments also underscore potential advantages of using code to evaluate effective semantic diversity. For code, our validity criteria (quality threshold) required both syntactic correctness and successful execution of all test cases without runtime errors; in contrast, for natural language, LLM-judge scores are less interpretable and may be susceptible to reward hacking. Additionally, code execution for diversity and quality assessment completes relatively quickly and incurs minimal cost, whereas LLM-judge evaluation can be computationally and financially expensive, necessitating careful selection of sample sizes to control API usage. Finally, in an environment where increasingly powerful models are being developed, using code to evaluate the diversity–quality trade-off may offer distinct advantages over relying on weaker models to evaluate potentially stronger models.
Comparison Validity (Quality) Soft Effective Semantic Diversity Hard Effective Semantic Diversity W (p) ES (d) Winner W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. <0.001<0.001 1.69 Inst. <0.001<0.001 0.80 Inst. <0.001<0.001 1.67 Inst. Small vs. Lg. <0.001<0.001 0.37 Lg. 0.004 0.47 Lg. 0.020 0.41 Lg. SFT vs. Pref. <0.001<0.001 2.13 Pref. 0.010 0.62 Pref. 0.050 0.58 Pref. SFT vs. DPO <0.001<0.001 1.33 DPO 0.092 0.42 DPO 0.176 0.39 DPO SFT vs. RL 0.004 5.80 RL 0.074 0.99 RL 0.203 0.83 RL DPO vs. RL 0.008 0.40 RL 0.129 -0.45 DPO 0.203 -0.29 DPO
Comparison Lexical Diversity Raw Neural Diversity W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. <0.001<0.001 1.30 Inst. <0.001<0.001 -3.21 Base Small vs. Lg. 0.202 -0.13 Lg. 0.010 -0.20 Small SFT vs. Pref. <0.001<0.001 0.95 Pref. <0.001<0.001 -2.38 SFT SFT vs. DPO <0.001<0.001 0.90 DPO <0.001<0.001 -1.51 SFT SFT vs. RL 0.004 1.70 RL 0.004 -5.57 SFT DPO vs. RL 0.570 -0.15 DPO 0.098 -0.37 DPO
Table 8: Creative Writing Model Comparison Results.
Comparison Validity (Quality) Soft Effective Semantic Diversity Hard Effective Semantic Diversity W (p) ES (d) Winner W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. <0.001<0.001 1.93 Inst. <0.001<0.001 0.86 Inst. <0.001<0.001 1.69 Inst. Small vs. Lg. 0.248 0.12 Lg. 0.026 0.54 Lg. 0.044 0.37 Lg. SFT vs. Pref. <0.001<0.001 1.55 Pref. 0.033 0.49 Pref. 0.032 0.41 Pref. SFT vs. DPO <0.001<0.001 1.07 DPO 0.213 0.26 DPO 0.266 0.22 DPO SFT vs. RL 0.008 2.40 RL 0.129 1.07 RL 0.129 1.14 RL DPO vs. RL 0.020 0.22 RL 0.263 0.10 RL 0.301 0.12 RL
Comparison Lexical Diversity Raw Neural Diversity W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. 0.033 0.16 Inst. <0.001<0.001 -3.07 Base Small vs. Lg. 0.010 -0.33 Small 0.594 -0.00 Small SFT vs. Pref. 0.919 -0.21 SFT <0.001<0.001 -1.90 SFT SFT vs. DPO 0.970 -0.25 SFT <0.001<0.001 -1.43 SFT SFT vs. RL 0.910 -0.14 SFT 0.004 -3.11 SFT DPO vs. RL 0.426 0.21 RL 0.359 -0.20 DPO
Table 9: Argumentative Writing Model Comparison Results.
Comparison Validity (Quality) Soft Effective Semantic Diversity Hard Effective Semantic Diversity W (p) ES (d) Winner W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. <0.001<0.001 1.27 Inst. <0.001<0.001 1.05 Inst. 0.002 0.86 Inst. Small vs. Lg. 0.859 -0.32 Small 0.645 -0.34 Small 0.546 -0.23 Small SFT vs. Pref. 0.013 0.37 Pref. 0.137 0.15 Pref. 0.919 -0.02 SFT SFT vs. DPO 0.339 0.14 DPO 0.970 0.01 DPO 0.151 -0.10 SFT SFT vs. RL 0.020 0.72 RL 0.039 0.34 RL 0.426 0.09 RL DPO vs. RL 0.055 0.24 RL 0.164 0.13 RL 0.203 0.12 RL
Comparison Lexical Diversity Raw Neural Diversity W (p) ES (d) Winner W (p) ES (d) Winner Base vs. Inst. <0.001<0.001 1.20 Inst. <0.001<0.001 -3.16 Base Small vs. Lg. 0.036 -0.58 Small <0.001<0.001 -0.48 Small SFT vs. Pref. 0.759 -0.13 SFT <0.001<0.001 -1.64 SFT SFT vs. DPO 0.176 -0.31 SFT <0.001<0.001 -1.32 SFT SFT vs. RL 0.301 0.16 RL 0.008 -2.19 SFT DPO vs. RL 0.027 0.45 RL 0.734 0.36 RL
Table 10: Brainstorming / Creative Assistance Model Comparison Results.